what is the solution of differential equation dy/dx=(y2+x2+2xy+y+x)/(y2+x2+2xy+2y+2x) ?

Grade:12

1 Answers

Abhishek Singh
9 years ago
(y2+x2+2xy+y+x)/(y2+x2+2xy+2y+2x)= ((y+x)2+(y+x))/((y+x)2+2(y+x))
                                                           = (x+y+1)/(x+y+2)
 
we can now solve the diff. eqn by substituting x+y=t